OpenStudy (anonymous):

A spring exerts a force of 5 N when stretc hed 1 m beyond its natural length of 2.4 m. How much work is done to stretch the spring from its natural length to a length of 4.2 m? a) find the spring constant b) how much work is required to stretch the spring from its natural length to a length of 4.2m?

OpenStudy (anonymous):

Use force = k * extenstion to find k where k is the spring constant

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

use the formula force = - k * x then u can find spring constant and the other required things

OpenStudy (shubhamsrg):

for the 2nd part, use W = -U = 1/2 kx^2

OpenStudy (shubhamsrg):

x= (4.2 -2.4)

OpenStudy (anonymous):

thank you! ill do that now